A train slow down from 200km/h With a uniform retardation of 10m/s². How long will it takes to reach 180km/h and what is the total distance covereds?


1
Expert's answer
2022-02-15T13:02:07-0500
"a=-10m\/s^2"

"a=v'=>v=v_0+at"

"v=v_0+at=s'=>s=s_0+v_0t+\dfrac{at^2}{2}"

Given "v_0=200km/h=\dfrac{200}{3.6}m/s"

"v_1=180km/h=\dfrac{180}{3.6}m/s"


"v_1=v_0+at_1"

"t_1=\dfrac{v_1-v_0}{a}=\dfrac{\dfrac{180}{3.6}m/s-\dfrac{200}{3.6}m/s}{-10m/s^2}="

"=\dfrac{5}{9}s\approx0.556 s"

If "s_0=0"


"s_1=\dfrac{200}{3.6}m/s(\dfrac{5}{9}s)+\dfrac{-10m/s^2(\dfrac{5}{9}s)^2}{2}=29.32m"

It takes "\dfrac{5}{9}s\approx0.556 s" to reach 180km/h. The total distance covered is "29.32m."